John is 2 years older than bill . If the sum of their ages is 36 , how old is john ?
Answer by ikleyn(52781)

.
John is 2 years older than bill . If the sum of their ages is 36 , how old is john ?
~~~~~~~~~~~~~~~~~~~~~~~~~~

J - B = 2,
J + B = 36.

Add the two equations.

2J = 2 + 36 = 38.

J = 38/2 = 19.

Answer.  John is 19 years old.